If the production function of the firm is Q = 50K 0.5L0.5   

 a) how will you derive the marginal product?

 b) what is the marginal value product with respect to L and with respect to K?

 c) how will the marginal rate of technical substitution look like for this production function?



1
Expert's answer
2021-01-29T20:00:45-0500

a) We can find the marginal product by partially differentiating the production function "Q" with respect to "L":


"MP_L=\dfrac{\partial Q}{\partial L}."

b) The marginal value product with respect to "L":


"MP_L=\dfrac{\partial}{\partial L}(50K^{0.5}L^{0.5})=\dfrac{25K^{0.5}}{L^{0.5}}."

The marginal value product with respect to "K":


"MP_K=\dfrac{\partial}{\partial K}(50K^{0.5}L^{0.5})=\dfrac{25L^{0.5}}{K^{0.5}}."

c) By the definition of the marginal rate of technical substitution, we have:


"MRTS_L^K=\dfrac{dL}{dK}=\dfrac{MP_L}{MP_K},""MRTS_L^K=\dfrac{\dfrac{25K^{0.5}}{L^{0.5}}}{\dfrac{25L^{0.5}}{K^{0.5}}}=\dfrac{K}{L}."